Given the demand for a product as Qd = 100 - 5P and the supply is given as Qs = -20 + 3p. You are told equilibrium is obtained at the point where Qd = Qs. The equilibrium price for the product is


1
Expert's answer
2021-06-07T10:36:01-0400

Equilibrium price and output

Equilibrium = Quantity Demanded (Qd) = Quantity Supplied (Qs)

Equilibrium price

"(Qd = 100-5P) = (Qs = -20 + 3P)"

"3P + 5P = 100 + 20"

"8P = 120"

"P= 120\/8"

"P = 15"